Half life of a certain radioactive substance is 69.3 days. Its disintegration constant is:

Grade:12

2 Answers

Adarsh
6 years ago
Dear student,
half life for 1st order reaction = 0.693/k
so   69.3=0.693/k    => k= 100.
Thanks and Regards 
Adarsh
 
anuragmishra
4 years ago
Half life for first order reaction= 0.693 / K
69.3 = 0.693/K
K= 0.01day-1
Here K is decay constant
Thanks and Regards 
Anurag Misra